OpenStudy (anonymous):

the total bill for repairing Steve's television was $25. the repair shop charged $25 an hour plus $19 for parts. how many hours of labor did it take to repair Steve's television? be sure to round the answer to the nearest tenth of an hour.

OpenStudy (anonymous):

0.24 hours

OpenStudy (anonymous):

how is it .24 hours

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

Total Charges = $ 25

OpenStudy (anonymous):

right or wrong

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Cost of Spares = $ 19

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so hour cost will be $(25-19) = $ 6

OpenStudy (anonymous):

so in one hour charges are $ 25 so how many hours will cost $ 6 \[\frac{6}{25} = 0.24 \approx 24%\]

OpenStudy (anonymous):

24 %

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

can u help me again

OpenStudy (anonymous):

the total bill for repairing Steve's television was $25. the repair shop charged $1 25 an hour plus $19 for parts. how many hours of labor did it take to repair Steve's television? be sure to round the answer to the nearest tenth of an hour.

OpenStudy (anonymous):

the total bill for repairing Steve's television was $125. the repair shop charged $1 25 an hour plus $19 for parts. how many hours of labor did it take to repair Steve's television? be sure to round the answer to the nearest tenth of an hour.

OpenStudy (anonymous):

same type of question u try it once by urself

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer 48 hours

OpenStudy (anonymous):

no how u calculated it tell me

OpenStudy (anonymous):

ok so total cost is $125 and the parts are $19 and one hour is $25 dollars so (25-19)= $6 and $^ divided by 125 is 0.048

OpenStudy (anonymous):

no total cost minus cost of parts = 125 - 19 = $ 106 cost per hour = $ 25 therefore $100 = 4 hours $ 100 + $6 = 4 hours + 0.24 hours = 4.24 hrs

OpenStudy (anonymous):

oo thank you